### v1

# Beyond Empathy: Integrating Diagnostic and Therapeutic Reasoning with Large Language Models for Mental Health Counseling

Abstract
Large language models (LLMs) hold significant potential for mental health support, capable of generating empathetic responses and simulating therapeutic conversations. However, existing LLM-based approaches often lack the clinical grounding necessary for real-world psychological counseling, particularly in explicit diagnostic reasoning aligned with standards like the DSM/ICD and incorporating diverse therapeutic modalities beyond basic empathy or single strategies. To address these critical limitations, we propose PsyLLM, the first large language model designed to systematically integrate both diagnostic and therapeutic reasoning for mental health counseling. To develop the PsyLLM, we propose a novel automated data synthesis pipeline. This pipeline processes real-world mental health posts, generates multi-turn dialogue structures, and leverages LLMs guided by international diagnostic standards (e.g., DSM/ICD) and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic) to simulate detailed clinical reasoning processes. Rigorous multi-dimensional filtering ensures the generation of high-quality, clinically aligned dialogue data. In addition, we introduce a new benchmark and evaluation protocol, assessing counseling quality across four key dimensions: comprehensiveness, professionalism, authenticity, and safety. Our experiments demonstrate that PsyLLM significantly outperforms state-of-the-art baseline models on this benchmark.
Beyond Empathy: Integrating Diagnostic and Therapeutic Reasoning with Large Language Models for Mental Health Counseling
He Hu1,2111Equal Contribution., Yucheng Zhou3111Equal Contribution., Juzheng Si4, Qianning Wang5, Hengheng Zhang2 Fuji Ren6, Fei Ma2222Corresponding Authors., Laizhong Cui1,2222Corresponding Authors. 1College of Computer Science and Software Engineering, Shenzhen University 2Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) 3SKL-IOTSC, CIS, University of Macau, 4Shandong University 5Auckland University of Technology, 6University of Electronic Science and Technology of China huhe@gml.ac.cn, yucheng.zhou@connect.um.edu.mo, mafei@gml.ac.cn, cuilz@szu.edu.cn
1 Introduction
Large language models (LLMs) have shown promising capabilities in mental health support tasks, such as generating empathetic responses and simulating therapeutic conversations Hegde and Jayalath (2025); Wang et al. (2024). Recent studies have explored incorporating emotion recognition and elements of cognitive behavioral therapy (CBT) to improve interpretability and control Xu et al. (2025); Zhang et al. (2024a). However, these systems often lack grounded reasoning and fail to reflect the diagnostic rigor required in real-world psychological counseling Na et al. (2025).
| Dataset | Turn Type | Reasoning? | Therapy? | Diagnostic? | Dialogues |
| SweetieChat Ye et al. (2025) | Multi-turn | ✗ | ✗ | ✗ | 3,757 |
| ESD Zhang et al. (2024b) | Multi-turn | ✗ | ✗ | ✗ | 2,900 |
| CPsyCounD Zhang et al. (2024a) | Multi-turn | ✗ | ✓ | ✗ | 3,134 |
| ChatCounselor Liu et al. (2023a) | Single-turn | ✗ | ✗ | ✗ | 8,187 |
| OpenR1-Psy (Ours) | Single + Multi-turn | ✓ | ✓ | ✓ | 19,302 |
A clinically grounded counseling system must account for both diagnostic rigor and therapeutic diversity. However, current LLM-based approaches fall short in the following two critical aspects: (1) Lack of deep reasoning on explicit diagnosis. In clinical practice, mental health professionals rely on international diagnostic standards such as the Diagnostic and Statistical Manual of Mental Disorders (DSM) and the International Classification of Diseases (ICD) to assess symptom patterns, determine diagnoses, and guide treatment planning Regier et al. (2013); Harrison et al. (2021). These standards ensure consistency, explainability, and legitimacy, especially in complex or comorbid cases. However, current counseling LLMs do not reason explicitly according to such frameworks, limiting their reliability in clinical decision support or realistic therapy simulation. (2) Neglect of therapeutic diversity. Not all patients benefit from the same therapeutic approach Gryesten et al. (2024); Castonguay et al. (2015); Beutler et al. (2006). Depending on the individual’s context, symptoms, and preferences, therapists may adopt different modalities such as cognitive behavioral therapy (CBT), acceptance and commitment therapy (ACT), psychodynamic therapy, or humanistic approaches Rowan (2006); Beck (1979); Hayes and Pierson (2005). This diversity is essential for effective care, but current models rarely incorporate multiple therapeutic perspectives or adapt their strategies accordingly.
To address existing limitations, we propose the first large language model that systematically integrates both diagnostic and therapeutic reasoning for mental health counseling, as shown in Fig. 1. Specifically, the model is capable of autonomously applying diagnostic criteria and various psychotherapeutic frameworks during counseling interactions. As shown in Table 1, we introduce a novel data synthesis pipeline that: (1) automatically parses real-world mental health posts; (2) generates simplified multi-turn diagnostic dialogues; (3) guides LLMs to simulate the reasoning processes of mental health professionals, incorporating step-by-step thinking aligned with DSM/ICD diagnostic logic and therapeutic considerations drawn from multiple evidence-based treatments; and (4) applies multi-dimensional filtering—including Incomplete Thinking, Context Incoherence, Mismatch Response, and No Framework—to ensure data quality. The resulting high-quality filtered dataset is used to fine-tune the LLM, enabling it to deliver counseling responses that are not only accurate and therapeutically grounded but also contextually appropriate and aligned with clinical reasoning processes.
In our experiments, we construct a benchmark to evaluate the quality of counseling responses along four key dimensions: comprehensiveness, professionalism, authenticity, and safety. Furthermore, we analyze the training data used for counseling responses, revealing that while increased model scale improves performance, the importance of data quantity is highly dependent on its quality. Our main contributions are as follows:
We propose PsyLLM, the first LLM that systematically integrates both diagnostic and therapeutic reasoning for mental health counseling.
We develop an automated data synthesis pipeline that generates high-quality, multi-turn dialogues reflecting diagnostic and therapeutic reasoning.
We incorporate international diagnostic standards (e.g., DSM/ICD) and diverse therapeutic strategies (e.g., CBT, ACT, psychodynamic) to guide our model’s deep reasoning.
We introduce a new benchmark and evaluation protocol tailored for counseling quality, and conduct a comprehensive analysis of model performance and data characteristics.
2 Related Work
2.1 Mental Support Chatbots
The increasing demand for accessible mental health services has spurred the development of AI-driven dialogue systems offering emotional support Ge et al. (2025). Early efforts, such as ChatCounselor Liu et al. (2023b) and PsyQA Sun et al. (2021), focused on single-turn, empathy-oriented responses, lacking therapeutic modeling or conversational depth. Subsequent work addressed multi-turn interaction: SMILE Qiu et al. (2024a) used ChatGPT to expand single replies into coherent dialogues, while SoulChat Chen et al. (2023) fine-tuned LLMs for better emotional adaptation across turns. SweetieChat Ye et al. (2025) introduced strategy-guided role-play to simulate diverse emotional scenarios, and CPsyCoun Zhang et al. (2024a) annotated dialogues with therapeutic types (e.g., CBT, humanistic), though it did not model underlying reasoning. To improve interpretability, ESCoT Zhang et al. (2024b) proposed a strategy-driven chain-of-thought process—emotion recognition, appraisal, and justification—mirroring human counseling. MentaLLaMA Yang et al. (2024) combined psychological insight with clinical condition prediction and rationales in social media settings. At the clinical edge, PsyGUARD Qiu et al. (2024b) addressed suicide risk detection and safety-aware response generation in multi-turn counseling.
2.2 Deep Reasoning in LLMs
Chain-of-Thought (CoT) prompting Wei et al. (2022) revealed LLMs’ capacity for multi-step inference, significantly improving symbolic reasoning. This has since extended to decision-making, multimodal understanding, and emotional inference. Dual-system models Dou et al. (2025) and modality-aware reinforcement learning Zhao et al. (2025) enhance adaptability and emotion recognition. In math reasoning, Wang et al. (2025) aligns vision and code for structured geometry tasks, outperforming GPT-4o. Dialogue reasoning introduces challenges in coherence and emotional depth; Shu et al. (2025) reframes CoT as agent dialogue, while Xie et al. (2025) fuses CoT with emotional cues from speech. Psychological reasoning is supported by CoT-modeled questionnaires Yang et al. (2023) and graph-based therapist–AI collaboration tools Chen et al. (2025). The full version of the related work can be found in Appendix A.
3 OpenR1-Psy Dataset Construction
The construction of the OpenR1-Psy dataset follows a systematic pipeline (Fig. 2), integrating data collection with LLM-powered parsing, therapeutic reasoning, and rigorous validation. We will detail each key component of dataset creation process.
3.1 Data Collection
To ensure both clinical realism and contextual diversity, our dataset sources are primarily divided into two parts: (1) The Reddit posts we used were pre-filtered and sourced from three publicly available corpora in mental health-related communities: Identifying-depression Pirina and Çöltekin (2018), Dreaddit Turcan and McKeown (2019), and LRF Garg et al. (2023). (2) Authentic patient utterances are derived from two high-quality psychological counseling datasets, ChatCounselor Liu et al. (2023b) and CPsyCoun Zhang et al. (2024a), which are based on real-world clinical cases and authentic patient-doctor dialogues, and refined as multi-turn dialogues between patients and LLM counselors by ChatGPT OpenAI (2023).
Rather than relying on isolated sentences, the post data we used is long-form posts that provide rich background information, including the poster’s personal context, psychological state, and social environment. These posts span diverse life situations, such as family dynamics, friendship challenges, interpersonal conflict, psychiatric symptoms, and financial stressors. In total, we collected 10,097 posts from these sources. After removing duplicates based on exact textual matching, we retained 8,752 unique posts.
3.2 Data Parsing and Interaction Planning
To convert raw Reddit posts into structured inputs for guiding dialogue generation, we implement a parsing and planning process inspired by psychological counseling techniques. This process uses a large language model (LLM), guided by a framework that simulates a counselor’s initial assessment and planning upon receiving a client’s self-report. For each Reddit post, the LLM performs the following analyses: (1) Emotion Assessment: Identify the user’s primary emotions, their intensity, and any emotional nuances. This step parallels how a therapist assesses affect. The output captures the key emotional states expressed. (2) Round Setting: Estimate the number of dialogue turns (1 to 3) based on the complexity of the issue. The dialogue progresses from surface-level emotions to deeper needs or vulnerabilities, reflecting typical therapeutic stages. (3) Theme Definition: Define the therapeutic focus for each round. These themes guide the conversation’s goals, ensuring each round builds on the previous one.
3.3 Question Extraction from Dialogue
Following the initial parsing and interaction planning, we prepare inputs for the deep reasoning process by creating a collection of patient utterances that will serve as prompts for counselor response generation. Let denote this collection of patient utterances. This is achieved through two mechanisms: (1) For the processed Reddit posts, we simulate a simplified initial empathetic multi-turn exchange using an LLM. Based on the planned dialogue structure (incorporating emotion assessment , round setting , and theme definition ), the LLM first generates a hypothetical empathetic counselor response to the post, followed by a plausible patient utterance in reaction to this turn. This patient’s utterance, representing a question or continuation of the user’s expression, is extracted as a prompt for the subsequent deep reasoning step. This effectively transforms static post data into dynamic dialogue starters, contributing to . (2) We incorporate data from existing high-quality psychological counseling datasets, specifically ChatCounselor Liu et al. (2023b) and CPsyCoun Zhang et al. (2024a). These datasets contain multi-turn patient-counselor dialogues. We extract only the patient utterances from these datasets, also contributing to .
The collection of patients’ turns from both sources forms the input pool for the subsequent generation step. This approach is central to our methodology: instead of utilizing pre-existing counselor responses, the collected patient utterances serve as prompts that guide the LLM to first generate a deep reasoning trace, which is then used to construct the final counselor response.
3.4 Response Generation with Diagnostic Standard & Therapeutic Strategy
Given a patient utterance (and potentially preceding dialogue context ), our system generates a clinically grounded counselor response . Unlike methods that directly map to and might offer post-hoc explanations, our approach explicitly models the internal reasoning process concurrently with generating the response. The LLM produces outputs that include both the reasoning trace and the counselor response in a single generation step, systematically guided by key principles of clinical practice: international diagnostic standards and diverse therapeutic strategies.
The integration of these elements is crucial for effective and responsible mental health support. International diagnostic standards, such as the Diagnostic and Statistical Manual of Mental Disorders (DSM) Regier et al. (2013) and the International Classification of Diseases (ICD) Harrison et al. (2021), provide a structured framework for understanding symptom patterns, recognizing potential underlying conditions, and ensuring a degree of diagnostic rigor that is fundamental in clinical assessment. Diverse therapeutic strategies, including but not limited to Cognitive Behavioral Therapy (CBT) Beck (1979), Acceptance and Commitment Therapy (ACT) Hayes and Pierson (2005), and psychodynamic approaches Rowan (2006), offer varied theoretical lenses and practical techniques to address the patient’s issues in a personalized and theoretically informed manner. Relying on a single strategy can limit applicability, as different patients respond better to different modalities.
Our system leverages a deep reasoning LLM to simulate this multi-faceted clinical thought process. Given and context , the LLM generates a structured output containing both the reasoning trace and the counselor utterance by considering relevant diagnostic perspectives (informed by , a representation of diagnostic knowledge) and evaluating applicable therapeutic strategies (informed by , a representation of therapeutic knowledge). This combined generation process can be conceptually represented as:
Here, and act as guiding frameworks that the LLM utilizes during the generation process. The model is specifically trained to first articulate the simulated clinical assessment and therapeutic rationale in , and then formulate the final counselor response conditioned on , , and the generated within this single output structure.
This joint generation of and ensures that the generated counseling response is not merely empathetic or plausible but is also demonstrably grounded in the clinical reasoning captured by . The reasoning is explicitly informed by diagnostic standards and therapeutic strategies , providing a level of interpretability and clinical legitimacy that is often lacking in end-to-end generative models. This aligns the LLM’s behavior more closely with the structured thought process of a human mental health professional, where the therapeutic intervention (the response) is directly derived from and justified by the clinical assessment and theoretical formulation (the reasoning).
3.5 Systematical Dialogue Validation
After the generation process, synthesized dialogue turns are constructed by combining patient utterances with their corresponding generated reasoning traces and counselor responses . To ensure the clinical relevance, logical consistency, and overall quality of this data for fine-tuning PsyLLM, we implement a rigorous validation process. This validation employs a multi-dimensional filtering mechanism applied to each generated turn based on and context . A sample is kept if it satisfies the following boolean condition :
where the boolean variables represent the evaluation along four key dimensions: , Complete Thinking Detection, which assesses the completeness and logical flow of the reasoning trace ; , Context Coherence Evaluation, which evaluates the consistency of with the preceding dialogue context and the user prompt ; , Mismatch Response Checking, which verifies the alignment between the reasoning trace and the counselor response ; and , Reasoning Framework Checking, which confirms whether adheres to the intended guidance from international diagnostic standards () and therapeutic strategies ().
Through this process, low-quality generated segments are filtered out, resulting in a high-quality dataset used for training. This systematic quality control is essential for building a reliable and clinically informed mental health counseling system.
3.6 Dataset Statistics and Analysis
| Category | Train | Test | All |
| # Dialogues | 18,852 | 450 | 19302 |
| # Utterances | 47,723 | 1651 | 49374 |
| Avg. Turn per Dialogue | 2.53 | 3.67 | 2.56 |
| Avg. Patient Utterance Length | 224.18 | 171.84 | 222.43 |
| Avg. Counselor Utterance Length | 520.11 | 489.4 | 519.08 |
| Avg. Reasoning Utterance Length | 1592.33 | 1627.36 | 1593.5 |
We analyze the constructed OpenR1-Psy dataset, presenting its key statistics and characteristics in Table 2 and Fig. 3. Table 2 shows the dataset statistics, comprising 19,302 dialogues (49,374 utterances) with a train/test split. The dialogues are multi-turn (avg. 2.56 turns). Patient utterances average 222 tokens, counselor responses 519 tokens, and notably, the reasoning trace averages 1593 tokens, highlighting its detailed nature. Fig. 3 presents key dataset distributions. (Left) shows psychotherapy approaches, dominated by Integrative (54.50%), Humanistic (25.15%), and CBT (17.16%), while including other diverse modalities. (Middle) illustrates scene categories, covering a broad range of topics like Emotion & Stress, Family Relationship, and Social Relationship. (Right) depicts severity levels: Moderate (48%) and Severe (41%) are most frequent, alongside Mild (10%) and Critical (1%), representing varied intensity.
3.7 PsyLLM Training
This section details the training process for PsyLLM, our large language model specialized for psychological counseling, leveraging the high-quality OpenR1-Psy dataset. The primary objective is to train the model to generate clinically informed counselor responses () grounded in explicit therapeutic reasoning (), given a patient’s utterance () and preceding dialogue context ().
We employ Supervised Fine-Tuning (SFT) on the OpenR1-Psy dataset. Each training instance uses the input sequence and the target output sequence , where the validated reasoning trace is concatenated before the corresponding counselor response . This structured format is crucial; it trains the model to first simulate the internal clinical reasoning process (generating ) and then formulate the response () conditioned on that reasoning and the input.
The training objective is to maximize the likelihood of generating the target sequence given by minimizing the cross-entropy loss over the target tokens. Formally, the model parameters are optimized by minimizing:
where is the number of training instances, is the length of the concatenated target sequence for instance , is the -th token, and are the preceding tokens in the target sequence.
This approach of jointly generating and , with preceding , ensures that the generated response is explicitly conditioned on the simulated clinical rationale . Unlike end-to-end models, this method enhances interpretability and clinical alignment by making the reasoning process traceable and integral to the response generation.
4 Experiments
4.1 Experimental Settings
Implementation Details.
To evaluate the effectiveness of the proposed OpenR1-Psy dataset in enhancing the reasoning and counseling capabilities of large language models, we fine-tune Qwen3-8B Yang et al. (2025) on the OpenR1-Psy dataset to obtain PsyLLM. The model is trained for 3 epochs using the AdamW optimizer with a learning rate of , standard parameters , and a weight decay of . A constant learning rate schedule was applied. We used a batch size of 8. Training was performed on 8 NVIDIA A100 (40G) GPUs.
Evaluation Metrics.
We employ a combination of automatic metrics and expert human review to rigorously assess the quality of the generated counselor responses and their accompanying reasoning traces. Our evaluation framework focuses on four key dimensions: comprehensiveness, assessing if the response addresses the core issues; professionalism, evaluating adherence to ethical guidelines and therapeutic principles; authenticity, ensuring the response feels natural and empathetic; and safety, verifying the absence of harmful or inappropriate content. These dimensions are applied to evaluate each generated single turn, which includes both the reasoning trace and the counselor utterance. We apply strict, reality-aligned criteria within these dimensions to ensure our evaluation reflects the standards expected in real-world therapeutic practice for individual responses.
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.740 | 1.830 | 1.430 | 0.990 | 0.574 |
| OpenR1-Psy | 1.630 (+120%) | 3.590 (+96%) | 2.660 (+86%) | 1.000 (+1%) | 0.900 (+57%) |
| CPsyCounD Zhang et al. (2024a) | 0.950 | 1.950 | 1.840 | 0.980 | 0.639 |
| OpenR1-Psy | 1.63 (+71%) | 3.59 (+84%) | 2.66 (+44%) | 1.00 (+2%) | 0.900 (+41%) |
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.015 | 0.826 | 0.332 | 0.469 | 0.202 |
| CPsyCounX Zhang et al. (2024a) | 0.005 | 0.492 | 0.180 | 0.358 | 0.137 |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 2.155 | 1.057 | 0.830 | 0.540 |
| GPT-4o Hurst et al. (2024) | 0.833 | 2.218 | 1.038 | 0.903 | 0.552 |
| PsyLLM | 0.904 | 2.210 | 1.440 | 0.945 | 0.607 |
| Method | Human | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| ChatCounselor Liu et al. (2023b) | 0.280 | 0.950 | 0.630 | 0.870 | 0.364 |
| CPsyCounX Zhang et al. (2024a) | 0.120 | 0.480 | 0.270 | 0.600 | 0.204 |
| DeepSeek-V3 DeepSeek-AI et al. (2024) | 0.880 | 1.880 | 1.250 | 0.810 | 0.507 |
| GPT-4o Hurst et al. (2024) | 0.850 | 2.170 | 1.270 | 0.920 | 0.541 |
| PsyLLM | 0.990 | 2.330 | 1.620 | 0.980 | 0.601 |
4.2 Synthetic Data Comparison
We conducted a synthetic data comparison to evaluate the quality of data generated by different methods. We randomly selected 150 patient cases from ChatCounselor Liu et al. (2023b) and 150 cases from CPsyCounD Zhang et al. (2024a), comparing their original counselor responses with those generated by our OpenR1-Psy method for the same inputs. Gemini-2.0-Flash Anil et al. (2023) evaluated these responses based on the four metrics: Comprehensiveness, Professionalism, Authenticity, and Safety (detailed prompts in Fig. 10).
Table 3 presents the evaluation results. OpenR1-Psy achieved substantial improvements over ChatCounselor, with gains of 120% in Comprehensiveness, 96% in Professionalism, 86% in Authenticity, and 1% in Safety, resulting in a 57% higher normalized average. Compared to CPsyCounD, OpenR1-Psy showed notable gains of 71% (Comprehensiveness), 84% (Professionalism), 44% (Authenticity), 2% (Safety), and 41% higher normalized average. These results confirm that the OpenR1-Psy generation method produces significantly higher quality data across key evaluation dimensions.
4.3 Performance Comparison
To evaluate PsyLLM’s effectiveness against state-of-the-art models, we conducted a turn-based dialogue evaluation on the OpenR1-Psy test set. We compared PsyLLM against baselines including ChatCounselor Liu et al. (2023b), CPsyCounX Zhang et al. (2024a), DeepSeek-V3 DeepSeek-AI et al. (2024), and GPT-4o OpenAI (2023). Evaluation utilized the four key metrics: Comprehensiveness, Professionalism, Authenticity, and Safety (with max scores outlined in Table 3). This process was automated using Gemini-2.0-Flash, guided by detailed prompts (refer to Fig. 11 in the Appendix). Table 4 shows the results. PsyLLM consistently demonstrates strong performance across all dimensions, achieving the highest scores in Comprehensiveness (0.904), Authenticity (1.440), and Safety (0.945). Although GPT-4o slightly edges out in Professionalism (2.218), PsyLLM’s score (2.210) is very close. Overall, PsyLLM attains the highest Normalized Average score (0.607), surpassing all baselines, including GPT-4o (0.552) and DeepSeek-V3 (0.540). These findings indicate that fine-tuning on the OpenR1-Psy dataset effectively enhances a language model’s capacity to generate clinically relevant and therapeutic counseling responses.
Human Evaluation.
To validate the automatic evaluation, we conducted a human evaluation. Five psychology graduate students assessed a randomly selected subset of dialogue turns from the test set, rating responses from PsyLLM and the baseline models (ChatCounselor Liu et al. (2023b), CPsyCounX Zhang et al. (2024a), DeepSeek-V3 DeepSeek-AI et al. (2024), GPT-4o Hurst et al. (2024)). Responses were evaluated blindly based on the four dimensions: Comprehensiveness, Professionalism, Authenticity, and Safety. As Table 5 shows, human evaluation confirms that PsyLLM consistently outperforms all baselines, achieving the highest scores across all four dimensions and the Normalized Average (0.601). It demonstrates PsyLLM’s superior ability to generate human-preferred and clinically counseling responses.
4.4 Analysis
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| Prompt w/o Clinical Frame | 1.075 | 2.994 | 2.079 | 0.979 | 0.740 |
| Prompt w/o Therapy Guidance | 0.803 | 2.871 | 2.053 | 0.985 | 0.697 |
| OpenR1-Psy | 1.083 | 3.277 | 2.299 | 0.988 | 0.778 |
| Method | Empathy | Clarity | Justification | Coherence | Structure | Normalized Avg |
| Human | 2.879 | 2.890 | 2.990 | 2.672 | 3.000 | 0.962 |
| Automatic Evaluation | 2.806 | 2.710 | 2.660 | 2.970 | 2.882 | 0.935 |
Impact of Model Scale.
To understand how the size of the underlying language model affects counseling performance after fine-tuning on OpenR1-Psy, we trained PsyLLM variants using different scales of the Qwen3 model series: Qwen3-1.7B, Qwen3-3B, and Qwen3-8B Yang et al. (2025). These models were fine-tuned using the same methodology and dataset as the main PsyLLM model. As shown in Figure 4 (Left), there is a clear trend of improving performance with increasing model size. The Qwen3-8B model achieves the highest overall Normalized Average score (0.757), indicating superior performance compared to the smaller models. The Qwen3-3B model performs better than the 1.7B version, achieving a Normalized Average of 0.738, while the smallest Qwen3-1.7B consistently shows the lowest overall performance with a Normalized Average of 0.675. These results indicate that larger base models, even when fine-tuned on the same dataset, possess stronger inherent capabilities that translate into better overall performance in generating nuanced, professional, and authentic psychological counseling responses.
Impact of Data Scale.
To investigate the influence of training data volume, we trained PsyLLM using varying proportions (20%, 40%, 60%, 80%, and 100%) of the OpenR1-Psy dataset, keeping the base model (Qwen3-8B) consistent. As shown in Figure 4 (Right), the impact of data scale on the Normalized Average is not strictly monotonic. Performance generally improves from 20% (0.718) up to 80% (0.746), with 80% achieving the highest Normalized Average. However, the Normalized Average drops significantly when using the full 100% of the training data (0.656). The Safety metric remained consistently high across all data scales (around 0.99). This indicates that simply using more data for model training does not always yield the best performance.
Effect of Diagnostic Standard & Therapeutic Strategy.
We conducted an ablation study on data generation by omitting either diagnostic guidance (“Prompt w/o Clinical Frame”) or therapeutic guidance (“Prompt w/o Therapy Guidance”), fine-tuning PsyLLM on these ablated datasets and the full OpenR1-Psy data. As seen in Table 6, training on the full OpenR1-Psy dataset yields the best Normalized Average (0.778). Removing therapeutic guidance results in a notable performance drop (0.697), especially in Comprehensiveness and Professionalism. Removing diagnostic guidance also reduces performance (0.740), though less significantly. It confirms that diagnostic standards and therapeutic strategies are critical.
Reasoning Process Evaluation.
A key aspect of our approach is the explicit generation of a reasoning trace. To evaluate its quality, we employed two sets of evaluators: five psychology graduate students (“Human”) and the Gemini model (“Automatic Evaluation”). Both groups scored the same set of PsyLLM-generated reasoning traces on five dimensions: Empathy, Clarity, Justification, Coherence, and Structure. Table 7 shows the average scores from each evaluator group. Human evaluators assigned slightly higher scores overall and on most dimensions, resulting in a higher Normalized Average (0.962 vs 0.935). Conversely, the automatic evaluator scored higher on Coherence. The high scores from both human and automatic evaluators indicate that PsyLLM can generate reasoning processes perceived as high-quality and comparable to human expert standards.
5 Conclusion
This paper addresses the critical limitations of existing LLMs for mental health support, specifically, their lack of explicit diagnostic reasoning (DSM/ICD aligned) and diverse therapeutic modality integration. We introduce PsyLLM, the first model designed to systematically integrate both, facilitated by a novel automated data synthesis pipeline creating the OpenR1-Psy dataset. This dataset incorporates explicit reasoning traces guided by diagnostic standards and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic), validated through multi-dimensional filtering. Our new benchmark, assessing counseling quality across comprehensiveness, professionalism, authenticity, and safety, demonstrates that PsyLLM significantly outperforms state-of-the-art baselines.
Limitations
PsyLLM performs well on short to medium-length dialogues but may struggle with long-term context tracking, complex emotional transitions, or maintaining nuanced therapeutic pacing across extended multi-turn interactions. Furthermore, due to limited computational resources and the lack of access to high-memory GPUs, PsyLLM is currently based on a 8B-sized model. Larger models with greater capacity for reasoning and emotional understanding have not been explored in this study.
Ethics Statements
Data Privacy
All data used in this study originate from publicly available datasets. The first portion consists of anonymous user-generated posts, in which users’ identities have been anonymized by the original data providers to ensure privacy (i.e., anonymous posting with no traceable user IDs). The second portion comprises an existing dialogue dataset, which underwent rule-based filtering, manual rewriting, and human proofreading to ensure the removal of any sensitive or privacy-related content.
Potential Risks of the Model
Although PsyLLM is optimized for therapeutic relevance, it has not been supervised or validated by licensed clinicians. Its responses should not be regarded as a replacement for professional mental health services. We conducted a thorough safety assessment of the training and evaluation data to ensure a high level of content security. However, the model fine-tuning process lacked human feedback, which means that certain responses may still carry potential risks of emotional harm, especially in sensitive scenarios.
Appendix A Related Work
A.1 Mental Support Chatbots
The growing demand for accessible and scalable mental health support has fueled the development of AI-driven dialogue systems capable of offering emotional and psychological assistance Ge et al. (2025). Early efforts in this area primarily centered on single-turn, empathy-oriented response generation. For instance, ChatCounselor Liu et al. (2023b) trained LLMs to produce emotionally appropriate responses evaluated across human-centric metrics. Similarly, PsyQA Sun et al. (2021) focused on generating supportive long-form answers in Chinese, but remained restricted to single-turn interactions without modeling therapeutic intent or multi-turn dynamics. To move beyond isolated responses, later work explored multi-turn emotional support. SMILE Qiu et al. (2024a) proposed a framework that expands single supportive utterances into coherent, multi-turn conversations using ChatGPT, enabling better contextual continuity. SoulChat Chen et al. (2023) extended this by fine-tuning LLMs on multi-turn empathetic dialogues, improving their ability to adapt across evolving emotional scenarios. Advancing further, SweetieChat Ye et al. (2025) introduced a strategy-guided role-playing framework to simulate diverse emotional situations while maintaining consistency across responses. In the Chinese counseling domain, CPsyCoun Zhang et al. (2024a) reconstructed multi-turn dialogues with therapeutic type annotations (e.g., CBT, humanistic), though it lacked modeling of the underlying reasoning behind those therapeutic strategies. To enhance transparency and interpretability, several works began to model the reasoning process itself. ESCoT Zhang et al. (2024b) presented an Emotion-Focused and Strategy-Driven Chain-of-Thought approach, making explicit the steps of emotion recognition, appraisal, and strategy justification in supportive responses—resembling human counseling logic. Similarly, MentaLLaMA Yang et al. (2024) focused on psychological insight in social media contexts by pairing clinical condition prediction with natural language rationales. At the clinical boundary, PsyGUARD Qiu et al. (2024b) tackled suicide risk detection and severity assessment in multi-turn counseling dialogues, emphasizing early triage and safety-aware response design.
A.2 Deep Reasoning in LLMs
Deep reasoning in LLMs has attracted increasing attention, particularly since the introduction of CoT prompting Wei et al. (2022), which encourages models to generate intermediate reasoning steps before producing final answers. This technique has shown notable improvements on arithmetic and symbolic tasks, revealing the latent ability of LLMs to perform multi-step inference. Recent work has extended deep reasoning capabilities beyond symbolic computation into more complex domains such as math reasoning, medical understanding, multimodal reasoning, and emotional understanding Zhou et al. (2025b, a); Hu et al. (2025); Zhou et al. (2024). For example, Dou et al. (2025) proposes a dual-system framework inspired by Kahneman’s theory of fast and slow thinking, integrating reinforcement learning and vision-language reasoning to support adaptive decision-making. Similarly, reinforcement learning has been leveraged to enhance both text-based reasoning and multimodal emotion recognition, with modality-specific contributions disentangled through verifiable rewards Zhao et al. (2025). In mathematical contexts, Wang et al. (2025) introduces a model that bridges vision and code by aligning image understanding with code generation, achieving superior performance in structured geometry problems compared to models like GPT-4o. As research pushes toward more interactive and human-centric applications, deep reasoning in dialogue systems has emerged as a challenging yet crucial frontier. Maintaining coherence, adaptability, and emotional sensitivity across multiple conversational turns introduces new demands. To address this, Shu et al. (2025) reformulates CoT prompting as a dialogue between reasoning agents and applies reinforcement learning to improve both coherence and solution diversity. In the context of empathetic conversations, Xie et al. (2025) proposes an LPE framework that combines CoT with emotional perception from spoken input, thereby guiding emotionally intelligent response generation. Moreover, psychological structures have been introduced to support deep reasoning in personality-related tasks: Yang et al. (2023) models psychological questionnaires as structured CoT chains for personality trait detection through multi-turn dialogue, while Chen et al. (2025) introduces a graph-based reasoning visualization system to facilitate collaboration between AI and therapists in therapeutic settings.
| Topic | Model | Comprehensiveness | Professionalism | Authenticity | Safety |
| Self-growth | ChatCounselor | 0.000 | 0.774 | 0.226 | 0.419 |
| CPsyCounX | 0.000 | 0.581 | 0.355 | 0.548 | |
| DeepSeek-V3 | 0.839 | 2.032 | 0.903 | 0.710 | |
| GPT-4o | 0.774 | 2.161 | 1.065 | 0.935 | |
| PsyLLM | 0.871 | 2.419 | 1.645 | 0.935 | |
| Emotion & Stress | ChatCounselor | 0.016 | 0.873 | 0.337 | 0.504 |
| CPsyCounX | 0.010 | 0.561 | 0.185 | 0.395 | |
| DeepSeek-V3 | 0.932 | 2.250 | 1.143 | 0.836 | |
| GPT-4o | 0.857 | 2.273 | 1.085 | 0.899 | |
| PsyLLM | 0.928 | 2.268 | 1.489 | 0.951 | |
| Education | ChatCounselor | 0.000 | 0.794 | 0.305 | 0.412 |
| CPsyCounX | 0.000 | 0.366 | 0.107 | 0.214 | |
| DeepSeek-V3 | 0.794 | 2.122 | 0.969 | 0.824 | |
| GPT-4o | 0.786 | 2.221 | 1.038 | 0.908 | |
| PsyLLM | 0.924 | 2.176 | 1.351 | 0.962 | |
| Love & Marriage | ChatCounselor | 0.018 | 0.788 | 0.343 | 0.472 |
| CPsyCounX | 0.005 | 0.374 | 0.152 | 0.326 | |
| DeepSeek-V3 | 0.833 | 1.992 | 0.962 | 0.833 | |
| GPT-4o | 0.838 | 2.162 | 0.982 | 0.919 | |
| PsyLLM | 0.881 | 2.141 | 1.386 | 0.952 | |
| Family Relationship | ChatCounselor | 0.021 | 0.785 | 0.343 | 0.453 |
| CPsyCounX | 0.003 | 0.467 | 0.183 | 0.349 | |
| DeepSeek-V3 | 0.869 | 2.114 | 1.017 | 0.827 | |
| GPT-4o | 0.830 | 2.163 | 0.993 | 0.889 | |
| PsyLLM | 0.889 | 2.135 | 1.398 | 0.941 | |
| Social Relationship | ChatCounselor | 0.017 | 0.922 | 0.353 | 0.457 |
| CPsyCounX | 0.000 | 0.655 | 0.250 | 0.362 | |
| DeepSeek-V3 | 0.897 | 2.362 | 1.181 | 0.810 | |
| GPT-4o | 0.776 | 2.345 | 1.095 | 0.888 | |
| PsyLLM | 0.888 | 2.276 | 1.500 | 0.905 | |
| Sex | ChatCounselor | 0.000 | 0.667 | 0.333 | 0.667 |
| CPsyCounX | 0.000 | 0.833 | 0.333 | 0.500 | |
| DeepSeek-V3 | 0.833 | 2.167 | 1.167 | 1.000 | |
| GPT-4o | 0.667 | 2.000 | 0.667 | 1.000 | |
| PsyLLM | 0.500 | 2.000 | 1.500 | 0.667 | |
| Career | ChatCounselor | 0.000 | 0.837 | 0.279 | 0.395 |
| CPsyCounX | 0.000 | 0.581 | 0.279 | 0.419 | |
| DeepSeek-V3 | 1.000 | 2.372 | 1.140 | 0.907 | |
| GPT-4o | 0.884 | 2.140 | 1.140 | 0.907 | |
| PsyLLM | 0.977 | 2.349 | 1.581 | 0.953 |
| Model | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| Qwen3-1.7B Yang et al. (2025) | 0.887 | 2.674 | 1.846 | 0.973 | 0.675 |
| Qwen3-3B Yang et al. (2025) | 0.969 | 3.101 | 2.125 | 0.985 | 0.738 |
| Qwen3-8B Yang et al. (2025) | 1.067 | 3.172 | 2.171 | 0.979 | 0.757 |
| Rate | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| 20% | 0.940 | 2.945 | 2.018 | 0.999 | 0.718 |
| 40% | 0.984 | 2.938 | 2.017 | 0.998 | 0.725 |
| 60% | 0.992 | 3.058 | 2.045 | 0.999 | 0.738 |
| 80% | 1.029 | 3.007 | 2.078 | 0.999 | 0.746 |
| 100% | 0.874 | 2.417 | 1.848 | 0.991 | 0.656 |
| Perspective | Criterion | Score | Subtotal |
| Comprehensiveness | 1.1 Accurate Empathy for Core Emotions: Does the response accurately identify and validate the client’s specific core emotions (e.g., anxiety, powerlessness, self-doubt) with precise language, going beyond generic phrases? | 1 | |
| 1.2 Insight into Underlying Needs & Pain Points: Does the response show insight into the client’s deeper, often unstated, emotional pain points or unmet needs (e.g., for safety, belonging, understanding) and gently guide exploration or affirm their validity? | 1 | 2 | |
| Professionalism | 2.1 Emotional Attunement & Validation: Does the counselor deeply empathize with and validate the client’s emotional experience, making the client feel "felt" and understood? | 1 | |
| 2.2 Natural Integration of Support: Is supportive guidance (reflection, affirmation, gentle questions) woven naturally into the conversation, without jargon or explicit theory-naming? | 1 | ||
| 2.3 Respectful & Non-Directive Style: Is the language consistently warm, respectful, and non-judgmental, avoiding commands or prescriptive "you should" statements? | 1 | 4 | |
| 2.4 Gentle Guidance & Client Agency: Are suggestions offered tentatively, encouraging client-led exploration and small steps, rather than providing direct solutions and respecting client autonomy? | 1 | ||
| Authenticity | 3.1 Conversational Fluency: Does the counselor communicate in a fluent, natural way, mirroring real therapeutic dialogue? | 1 | |
| 3.2 Relational Presence: Is there evidence of genuine emotional connection and relational presence with the client? | 1 | 3 | |
| 3.3 Pacing and Responsiveness: Does the counselor maintain appropriate pacing, leaving space for client expression, and respond in a natural, conversational manner to the client’s immediate input? | 1 | ||
| Safety | 4.1 Ethical Boundaries & Client Protection: Does the counselor avoid intrusive or leading language, respect privacy by not probing for unnecessary sensitive details, and prioritize client well-being and autonomy? | 0.5 | |
| 4.2 Fostering Psychological Safety: Does the counselor consistently convey unconditional respect and acceptance, creating an environment free of judgment or dismissal, where the client feels safe to be authentic? | 0.5 | 1 |